1.   Assume that the exam result of Data Science class fits a normal distribution. Furthermore, the mean exam result  is 82, and the standard deviation is 17.2. What is the percentage of students scoring 90 or more in the exam?


1
Expert's answer
2022-01-18T09:18:53-0500

Solution:

"\mu=82, \sigma=17.2
\\ X\sim N(\mu,\sigma)
\\ P(X\ge 90)=1-P(X \le 90)
\\=1-P(z\le \dfrac{90-82}{17.2})
\\=1-P(0.47)
\\=1-0.68082
\\=0.31918"